25.0 g of a molecular compound is dissolved in 50.0 g of water giving the resulting solution a boiling point of 104 ˚C. Calculate the molar mass of the solution. K
b
of water is 0.52 ˚C/m.
Multiple Choice
65.0 g/mol
75.4 g/mol
107.6 g/mol
122.3 g/mol
1
Expert's answer
2020-12-11T02:49:04-0500

ΔT = T(boiling solution) – T(water boiling)

ΔT = 104 – 100 = 4 ºC

"m = \frac{ΔT}{K_b} \\

m = \frac{4}{0.52} = 7.69 \;mol/kg \\

m = \frac{moles\;solute}{kg\;solvetnt}"

kg solvent = 0.05 kg

moles solvent "= m \times kg\;solvent"

moles solvent "= 7.69 \times 0.05 = 0.3845 \;mol"

"m_c = 25.0 \;g \\

n_c = 0.3845 \;mol \\

MM_c = \frac{25.0}{0.3845} = 65.02 \;g/mol"

Answer: 65.02 g/mol